Two objects are involved in a collision. The first object, of mass 12 kg slows from 8 m s-1 to zero on 0.5 s applying its force to the second object which has a mass of 10 kg. Calculate:
a) the acceleration to the first object.
b) the force applied to the second object.
c) the acceleration of the second object during the collision.
1
Expert's answer
2019-04-18T11:34:00-0400

a)


"a_1=\frac{8-0}{0.5}=16 \frac{m}{s^2}"

b)


"F=m_1a_1=(12)(16)=192N"

c)


"a_2=\frac{F}{m_2}=\frac{192}{10}=19.2 \frac{m}{s^2}"